# One More Glance with Sharp Eyes: Rethinking Lightweight Captioning as a Practical Visual Specialist

### v1

@@ 7 Related Work
have attracted considerable research attention due to their versatile applications, such as chat-bots [80]. Early approaches integrated contrastive image-language pretrained models [74] with powerful LLMs, enabling complex reasoning. The development of instruction-based datasets [43] and innovative training strategies [35; 86] have further accelerated progress, substantially improving MLLM performance and broadening their capabilities. Despite these achievements, most MLLMs heavily depend on large-parameter LLMs, making deployment on memory-constrained devices unfeasible. This limitation will likely restrict access for a significant portion of users worldwide.
@@
Recent advances in image captioning have improved training efficiency and descriptive fidelity. Approaches such as CaMEL [6] and SmallCap [58] emphasize minimizing trainable parameters by leveraging mean-teacher distillation and employing retrieval augmentation, while Tag2Text [25] and LoCCa [72] introduce novel mechanisms such as dedicated tagging and location-aware refinement to improve caption quality. Unlike existing approaches that focus on reducing trainable parameters, or rely on single-pass inference—potentially missing crucial details—our method prioritizes inference efficiency considering on-device operation and systematically addressing the limitations of single-pass generation.
Visual Blindness in VLMs.
@@
Humans often refine their writing through iterative review to enhance clarity and precision [50; 20]. Recent research has applied this refinement concept to LLMs, introducing techniques. For instance, Self-Refine [32; 59; 54] enabled models to autonomously critique and iteratively enhance their outputs. Unlike such approaches limited to the LLM domain, our method introduces refinement in a multimodal context, guided by both language and vision.
@@ 8 Conclusion & Broader Impact
We demonstrate the potential of small-scale specialists, highlighting their surprising effectiveness. As a next step, future research may explore iterative refinement strategies, integrating external tools (e.g., zoom or crop, as in GPT-4o [52]), and designing a unified multimodal connector for both the initial and secondary glance. More research questions that aim to enhance both captioning performance and efficiency in real-world applications are provided in Sec. D.6.
@@ Appendix A Summary of Our Work
    Factual image captioning does not significantly require the complex reasoning abilities of LLMs and can be handled by a lightweight language model.
@@
    Single-pass captioning can lead specialists to adopt broad attention, neglecting critical visual regions necessary.
@@
    CLIP’s final-layer features provide only coarse-grained visual information, limiting their suitability for fine-grained captioning tasks.
@@
  • Our novel insight: [Uncaptioned image] Encouraging attention to key regions via the initial caption, along with multi-level utilization of a single vision encoder, can be an effective and efficient strategy.
@@
    Unlike traditional works, we compared not only existing small specialist models but also large multimodal models, demonstrating their potential as strong visual specialists.
@@
    Industry: Our model can be easily applied in industrial scenarios such as assistance devices for the visually impaired or navigating robots. Our model runs successfully on actual edge devices like Jetson Nano.
@@
Appendix B Additional Experiments
B.1 When GPT meets ‘one more glance’
In this work, we suggest that allowing the model to take one more glance, guided by the initial caption and aligned visual features, offers an effective approach, as discussed in Note 4. We further explore how OpenAI’s GPT models, which have demonstrated remarkable performance, might be influenced by our refinement framework. While our framework consists of two components, taking one more glance and refining with sharp eyes through DeepLens, we apply only the former as modifying their internal architecture is not feasible. The results are shown in Fig. 10, where GPT-4-turbo is used. As a commercial model with a sufficiently large number of parameters, it often produces an accurate initial caption without noticeable errors, but we still observed occasional inaccuracies in its outputs. To explore whether refinement could help, we instructed the model to take one more glance and found that GPT can generate a refined caption by re-examining the given image. OpenAI’s GPT-o3 reflects an awareness of this idea through the introduction of the Thinking with Images [52]. However, this concept remains less explored in the research community, suggesting the need for further investigation.
B.2 When larger specialists meet ‘one more glance with sharp eyes’
Building on our lightweight specialist, we naturally ask: since our framework and module offer a scalable methodology, can they also be effective when applied to larger multimodal models? To investigate this, we scaled the language model from OPT-125M to OPT-1.3B and also utilized LLaMA-2-7B. Specifically, we fine-tuned these models on the ShareGPT and DCI training datasets to serve as captioning specialists (i.e., single pass). Next, we incorporated our Sharp-Eyed Refinement (i.e., SeR) framework with DeepLens and further fine-tuned them using the datasets described in Sec. 5. As shown in Tab. 5, our framework consistently improves caption accuracy across different models. In particular, we observe gains of 1.2 and 0.9 points in CAPT [18] evaluation for OPT-1.3B and LLaMA-2-7B, respectively. Additionally, the results indicate that using larger language models to build captioning specialists leads to improved performance, observed in prior LLM scaling studies [30, 79, 24]. Despite the gains observed with larger models such as OPT-1.3B and LLaMA-2-7B, we note that their considerable size (roughly 10x and 56x larger than OPT-125M, respectively) may hinder real-world applicability. It is therefore crucial to continue exploring captioning systems that balance effectiveness with efficiency.
@@ B.2 When larger specialists meet ‘one more glance with sharp eyes’
B.3 Evaluation on an Actual Edge Device
@@
Here, we demonstrate that lightweight captioners can be effectively deployed on edge devices such as RTX 3090, and Jetson Nano. We evaluate captioning performance on MS COCO and ShareGPT4V&DCI datasets, and additionally measure inference time, memory usage, and power consumption. All models were executed for 100 iterations with a batch size of 1. The results in Tab. 6 show that our method consistently performs well across devices, and notably, remains fully operational on setups where models like LLaVA-1.5 fail to run. These findings support the deployability of our framework in edge scenarios and encourage further exploration of lightweight captioning for real-world assistive technologies.
B.4 Comparison on different learning strategies
As writers frequently revise their drafts to improve clarity or tone, and programmers refactor code to enhance structure, we demonstrate that allowing multimodal models to refine their own outputs through our framework can be an effective strategy. In this subsection, we examine how our lightweight specialist performs under different learning strategies. The first strategy involves training with as much data as possible. Here, we train the model on a combined set of COCO, ShareGPT, DCI, and GLaMM to maximize data exposure (i.e., more data). The second strategy explores distillation using teacher-generated captions. Specifically, we train our model on captions produced by LLaVA-1.6-34B [44] (i.e., distillation). Our original strategy was straightforwardly fine-tuning the model exclusively on task-specific data (i.e., origin). The results are summarized in Tab. 7. For the ‘more data’ strategy, we observe limited performance, likely due to reduced task alignment caused by mixing heterogeneous datasets. In contrast, the ‘distillation’ strategy demonstrates that learning from a strong teacher can improve performance even in a single-pass setting. Notably, when combined with our refinement framework (SeR), it yields further gains, suggesting that distillation and SeR can be effectively integrated. Additionally, as noted in Sec. 5, the datasets we used share similarities with those in Direct Preference Optimization (DPO). Investigating the impact of training our model using reinforcement learning methods such as DPO could be a promising direction for future work.
@@ B.4 Comparison on different learning strategies
B.5 Efficacy of Multi-Level Features with Other Vision Encoders
We investigate whether our strategy of selecting multi-level features can generalize to vision encoders beyond CLIP [56], which is our original setup. As shown in Tab. 8, the approach consistently improves performance across multiple encoders. Specifically, both SigLIPv2 [68] and DINOv2 [53] exhibit performance gains when equipped with multi-level features, demonstrating that our method is not limited to CLIP-based models. Interestingly, CLIP with multi-level features outperforms the combination of CLIP and DINOv2 in several evaluation metrics, while maintaining greater parameter efficiency. On the other hand, DINOv2 alone yields relatively lower performance, likely due to weaker alignment with language features. For these experiments, we pair OPT-125M with each vision encoder and fine-tune the resulting models on the ShareGPT [11]&DCI [70] datasets. Subsequently, we apply our Sharp-Eyed Refinement framework using the selected multi-level features and perform second fine-tuning with the datasets described in Sec. 5 of the main paper.
@@ B.5 Efficacy of Multi-Level Features with Other Vision Encoders
B.6 Iterative Sharp-Eyed Refinement
To explore the potential of iterative refinement beyond a single pass, we conduct additional experiments using captioning specialists with larger model capacity. This decision is motivated by our initial hypothesis that the constrained capacity of small models may have hindered the effectiveness of multi-step refinement. As shown in Tab. 9, applying multiple refinement iterations to the smallest language model, OPT-125M, results in only limited improvements. In contrast, we observe that multi-step refinement leads to performance gains when used with higher-capacity captioners. These findings underscore the potential of scaling up refinement strategies, as similarly observed in recent LLM-based studies [50, 59, 81, 31]. We plan to further investigate this direction in future work.
For implementation, we construct larger captioning specialists based on two configurations: LLaMA-2-7B + CLIP-ViT-Large and OPT-1.3B + CLIP-ViT-Large. These models are initially fine-tuned on the ShareGPT and DCI training datasets via single-pass training. We then apply the Sharp-Eyed Refinement (SeR) framework with DeepLens in multiple iterations, followed by additional fine-tuning using the datasets described in Sec. 5.
@@ B.6 Iterative Sharp-Eyed Refinement
Appendix C DeepLens
C.1 Architectural details
We have shown that our refinement framework can improve caption quality. At the core of it lies DeepLens, which we hypothesize plays a critical role by aligning multi-level visual features with semantic cues from the initial caption. As illustrated in Fig. 11, DeepLens takes two new inputs: ① the previously generated caption, represented as token embeddings, and ② multi-layer features extracted from the vision encoder. Let $N$, $d$, and $m$ denote the number of visual tokens, the embedding dimension, and the number of selected layers, respectively. The selected features are concatenated along the channel dimension to retain hierarchical visual information. These fused inputs are then passed through Transformer layers following the BERT architecture [17]. The resulting output is provided to the language model, enabling refinement conditioned on a focused and semantically grounded visual representation.
Figure 11: Illustration of the DeepLens architecture, which integrates ① token embeddings from the initial caption and ② multi-layer visual features for refinement.
C.2 Ablation study
We conduct a series of ablation experiments to analyze the design choices behind DeepLens, as summarized in Tab. 10. The corresponding model variants are illustrated in Fig. 12. Our study spans three key aspects: connector selection, refinement configurations, and the layer indexing strategy for visual feature extraction. (I) We begin by comparing different multimodal connectors used in the initial caption generation phase. While prior works have explored a range of designs, such as Cross-Attention [58, 6], Q-Former [38], and Transformer-style modules, we find that the simple MLP connector from LLaVA [42] offers competitive performance. Notably, (b) and (f), which adopt multi-level features and a BERT-style Transformer block respectively, perform slightly better than the MLP-based connector. However, given the marginal gains, we retain the lightweight MLP structure for efficiency. (II) Next, we evaluate several connectors for the refinement phase (SeR). Among them, combining our base configuration (a) with structure (A), which utilizes both proposed inputs, yields strong performance. Although we also applied (A) to configuration (b), the improvement was marginal, so we did not adopt it as our baseline. (III) Finally, we investigate the impact of selecting different sets of layers from the ViT encoder. We test multiple combinations and find that using a diverse set of layers {13, 18, 23} results in the highest performance.
Figure 12: Architectural designs of the model variants used in the ablation study. Each configuration corresponds to a specific component or setting evaluated in Tab. 10
@@ C.2 Ablation study
Table 10: Ablation results evaluating connector types, refinement configurations, and ViT layer selections in DeepLens. Architectural variants are illustrated in Fig. 12.
Appendix D Discussions
D.1 Why previous small models fall short?
@@
In contrast to this trend, we analyze small captioning models and uncover the following: Aforementioned small captioning models typically adopt an architecture where visual features are injected via cross-attention. Unfortunately, when using a small language model, this design does not appear to be particularly beneficial. Empirically, this limitation is reflected in the performance of structure (c) in Fig. 12, which yields a relatively low score of 125.9 as shown in Tab. 10. Had the field not been so heavily influenced by the rise of LLMs, such limitations in small model architectures might have been identified and addressed earlier.
@@
We adopt the LLaVA architecture and inject visual features into the self-attention inputs, as shown in Fig. 12 (a). We show that this simple setup surprisingly delivers stronger performance (i.e., Note 1 in Sec. 3) and contributes to demonstrating that small models can be viable for practical applications. We hope this encourages a shift toward reducing dependence on LLMs in specific tasks like captioning, and sparks further interest in the development of lightweight yet effective models.
D.2 Limitations of Existing Multimodal Large Language Models

### v2

@@ 7 Related Work
have attracted considerable research attention due to their versatile applications, such as chat-bots wu2023multimodal . Early approaches integrated contrastive image-language pretrained models wang2022git with powerful LLMs, enabling complex reasoning. The development of instruction-based datasets llava and innovative training strategies li2022mplug ; ye2024mplug have further accelerated progress, substantially improving MLLM performance and broadening their capabilities. Despite these achievements, most MLLMs heavily depend on large-parameter LLMs, making deployment on memory-constrained devices unfeasible. This limitation will likely restrict access for a significant portion of users worldwide.
@@
Recent advances in image captioning have improved training efficiency and descriptive fidelity. Approaches such as CaMEL barraco2022camel and SmallCap ramos2023smallcap emphasize minimizing trainable parameters by leveraging mean-teacher distillation and employing retrieval augmentation, while Tag2Text huang2023tag2text and LoCCa wan2024locca introduce novel mechanisms such as dedicated tagging and location-aware refinement to improve caption quality. Unlike existing approaches that focus on reducing trainable parameters, or rely on single-pass inference—potentially missing crucial details—our method prioritizes inference efficiency considering on-device operation and systematically addressing the limitations of single-pass generation.
@@
Humans often refine their writing through iterative review to enhance clarity and precision selfrefine ; dou2024stepcoder . Recent research has applied this refinement concept to LLMs, introducing techniques. For instance, Self-Refine le2022coderl ; ranaldi2024self ; pan2023automatically enabled models to autonomously critique and iteratively enhance their outputs. Unlike such approaches limited to the LLM domain, our method introduces refinement in a multimodal context, guided by both language and vision.
@@ 8 Conclusion & Broader Impact
We demonstrate the potential of small-scale specialists, highlighting their surprising effectiveness. As a next step, future research may explore iterative refinement strategies, integrating external tools (e.g., zoom or crop, as in GPT-4o openai2024o3o4 ), and designing a unified multimodal connector for both the initial and secondary glance. More research questions that aim to enhance both captioning performance and efficiency in real-world applications are provided in Section E.6.
@@ Appendix A Summary of Our Work
    (Note1) Factual image captioning does not significantly require the complex reasoning abilities of LLMs and can be handled by a lightweight language model.
@@
    (Note2) Single-pass captioning can lead specialists to adopt broad attention, neglecting critical visual regions necessary.
@@
    (Note3) CLIP’s final-layer features provide only coarse-grained visual information, limiting their suitability for fine-grained captioning tasks.
@@
  • (Note4) Encouraging attention to key regions via the initial caption, along with multi-level utilization of a single vision encoder, can be an effective and efficient strategy.
@@
    We compared not only existing small captioning models but also large multimodal models, demonstrating the potential of our model as a strong visual specialist.
  3. 3.
    Through these extensive experiments, (Note5) we show that our proposed baseline and framework have the potential of lightweight captioning for on-device applications.
@@
    Industry: Our model can be easily applied in industrial scenarios such as assistive devices for the visually impaired or navigation robots. Our model runs successfully on actual edge devices like Jetson Nano.
@@
Appendix B Additional Related Work
B.1 Visual Blindness in VLMs.
Appendix C Additional Experiments
C.1 When GPT meets ‘one more glance’
In this work, we suggest that allowing the model to take one more glance, guided by the initial caption and aligned visual features, offers an effective approach, as discussed in Note 4. We further explore how OpenAI’s GPT models, which have demonstrated remarkable performance, might be influenced by our refinement framework. While our framework consists of two components, taking one more glance and refining with sharp eyes through DeepLens, we apply only the former as modifying their internal architecture is not feasible. The results are shown in Figure 10, where GPT-4-turbo is used. As a commercial model with a sufficiently large number of parameters, it often produces an accurate initial caption without noticeable errors, but we still observed occasional inaccuracies in its outputs. To explore whether refinement could help, we instructed the model to take one more glance and found that GPT can generate a refined caption by re-examining the given image. OpenAI’s GPT-o3 reflects an awareness of this idea through the introduction of the Thinking with Images [53]. However, this concept remains less explored in the research community, suggesting the need for further investigation.
C.2 When larger specialists meet ‘one more glance with sharp eyes’
Building on our lightweight specialist, we naturally ask: since our framework and module offer a scalable methodology, can they also be effective when applied to larger multimodal models? To investigate this, we scaled the language model from OPT-125M to OPT-1.3B and also utilized LLaMA-2-7B. Specifically, we fine-tuned these models on the ShareGPT and DCI training datasets to serve as captioning specialists (i.e., single pass). Next, we incorporated our Sharp-Eyed Refinement (i.e., SeR) framework with DeepLens and further fine-tuned them using the datasets described in Section 5. As shown in Table 5, our framework consistently improves caption accuracy across different models. In particular, we observe gains of 1.2 and 0.9 points in CAPT [18] evaluation for OPT-1.3B and LLaMA-2-7B, respectively. Additionally, the results indicate that using larger language models to build captioning specialists leads to improved performance, observed in prior LLM scaling studies [30, 79, 24]. Despite the gains observed with larger models such as OPT-1.3B and LLaMA-2-7B, we note that their considerable size (roughly 10x and 56x larger than OPT-125M, respectively) may hinder real-world applicability. It is therefore crucial to continue exploring captioning systems that balance effectiveness with efficiency.
@@ C.2 When larger specialists meet ‘one more glance with sharp eyes’
C.3 Evaluation on an Actual Edge Device
@@
Here, we demonstrate that lightweight captioners can be effectively deployed on edge devices such as RTX 3090 and Jetson Nano. We evaluate captioning performance on MS COCO and ShareGPT4V&DCI datasets, and additionally measure inference time, memory usage, and power consumption. All models were executed for 100 iterations with a batch size of 1. The results in Table 6 show that our method consistently performs well across devices, and notably, remains fully operational on setups where models like LLaVA-1.5 fail to run. These findings support the deployability of our framework in edge scenarios and encourage further exploration of lightweight captioning for real-world assistive technologies.
C.4 Comparison on different learning strategies
As writers frequently revise their drafts to improve clarity or tone, and programmers refactor code to enhance structure, we demonstrate that allowing multimodal models to refine their own outputs through our framework can be an effective strategy. In this subsection, we examine how our lightweight specialist performs under different learning strategies. The first strategy involves training with as much data as possible. Here, we train the model on a combined set of COCO, ShareGPT, DCI, and GLaMM to maximize data exposure (i.e., more data). The second strategy explores distillation using teacher-generated captions. Specifically, we train our model on captions produced by LLaVA-1.6-34B [45] (i.e., distillation). Our original strategy was straightforwardly fine-tuning the model exclusively on task-specific data (i.e., origin). The results are summarized in Table 7. For the ‘more data’ strategy, we observe limited performance, likely due to reduced task alignment caused by mixing heterogeneous datasets. In contrast, the ‘distillation’ strategy demonstrates that learning from a strong teacher can improve performance even in a single-pass setting. Notably, when combined with our refinement framework (SeR), it yields further gains, suggesting that distillation and SeR can be effectively integrated. Additionally, as noted in Section 5, the datasets we used share similarities with those in Direct Preference Optimization (DPO). Investigating the impact of training our model using reinforcement learning methods such as DPO could be a promising direction for future work.
@@ C.4 Comparison on different learning strategies
C.5 Efficacy of Multi-Level Features with Other Vision Encoders
We investigate whether our strategy of selecting multi-level features can generalize to vision encoders beyond CLIP [57], which is our original setup. As shown in Table 8, the approach consistently improves performance across multiple encoders. Specifically, both SigLIPv2 [68] and DINOv2 [54] exhibit performance gains when equipped with multi-level features, demonstrating that our method is not limited to CLIP-based models. Interestingly, CLIP with multi-level features outperforms the combination of CLIP and DINOv2 in several evaluation metrics, while maintaining greater parameter efficiency. On the other hand, DINOv2 alone yields relatively lower performance, likely due to weaker alignment with language features. For these experiments, we pair OPT-125M with each vision encoder and fine-tune the resulting models on the ShareGPT [11]&DCI [70] datasets. Subsequently, we apply our Sharp-Eyed Refinement framework using the selected multi-level features and perform second fine-tuning with the datasets described in Section 5 of the main paper.
@@ C.5 Efficacy of Multi-Level Features with Other Vision Encoders
C.6 Iterative Sharp-Eyed Refinement
To explore the potential of iterative refinement beyond a single pass, we conduct additional experiments using captioning specialists with larger model capacity. This decision is motivated by our initial hypothesis that the constrained capacity of small models may have hindered the effectiveness of multi-step refinement. As shown in Table 9, applying multiple refinement iterations to the smallest language model, OPT-125M, results in only limited improvements. In contrast, we observe that multi-step refinement leads to performance gains when used with higher-capacity captioners. These findings underscore the potential of scaling up refinement strategies, as similarly observed in recent LLM-based studies [51, 60, 82, 32]. We plan to further investigate this direction in future work.
For implementation, we construct larger captioning specialists based on two configurations: LLaMA-2-7B + CLIP-ViT-Large and OPT-1.3B + CLIP-ViT-Large. These models are initially fine-tuned on the ShareGPT and DCI training datasets via single-pass training. We then apply the Sharp-Eyed Refinement (SeR) framework with DeepLens in multiple iterations, followed by additional fine-tuning using the datasets described in Section 5.
@@ C.6 Iterative Sharp-Eyed Refinement
Appendix D DeepLens
D.1 Architectural details
We have shown that our refinement framework can improve caption quality. At the core of it lies DeepLens, which we hypothesize plays a critical role by aligning multi-level visual features with semantic cues from the initial caption. As illustrated in Figure 11, DeepLens takes two new inputs: ① the previously generated caption, represented as token embeddings, and ② multi-layer features extracted from the vision encoder. Let $N$, $d$, and $m$ denote the number of visual tokens, the embedding dimension, and the number of selected layers, respectively. The selected features are concatenated along the channel dimension to retain hierarchical visual information. These fused inputs are then passed through Transformer layers following the BERT architecture [17]. The resulting output is provided to the language model, enabling refinement conditioned on a focused and semantically grounded visual representation.
Figure 11: (i) Illustration of the DeepLens architecture. This is designed for integrating ① token embeddings from the initial caption and ② multi-layer visual features for refinement. (ii) and (iii) Ablation study of multimodal connector designs. Each configuration corresponds to evaluation in Table 10.
D.2 Ablation study
We conduct a series of ablation experiments to analyze the design choices behind DeepLens, as summarized in Table 10. The corresponding model variants are illustrated in Figure 11. Our study spans three key aspects: connector selection, refinement configurations, and the layer indexing strategy for visual feature extraction. (I) We begin by comparing different multimodal connectors used in the initial caption generation phase. While prior works have explored a range of designs, such as Cross-Attention [59, 6], Q-Former [39], and Transformer-style modules, we find that the simple MLP connector from LLaVA [43] offers competitive performance. Notably, (b) and (f), which adopt multi-level features and a BERT-style Transformer block respectively, perform slightly better than the MLP-based connector. However, given the marginal gains, we retain the lightweight MLP structure for efficiency. (II) Next, we evaluate several connectors for the refinement phase (SeR). Among them, combining our base configuration (a) with structure (A), which utilizes both proposed inputs, yields strong performance. Although we also applied (A) to configuration (b), the improvement was marginal, so we did not adopt it as our baseline. (III) Finally, we investigate the impact of selecting different sets of layers from the ViT encoder. We test multiple combinations and find that using a diverse set of layers {13, 18, 23} results in the highest performance.
@@ D.2 Ablation study
Table 10: Ablation results evaluating connector types, refinement configurations, and ViT layer selections in DeepLens. Architectural variants are illustrated in Figure 11 (b) and (c).
Appendix E Discussions
E.1 Why do previous small models fall short?
@@
In contrast to this trend, we analyze small captioning models and uncover the following: Aforementioned small captioning models typically adopt an architecture where visual features are injected via cross-attention. Unfortunately, when using a small language model, this design does not appear to be particularly beneficial. Empirically, this limitation is reflected in the performance of structure (c) in Figure 11, which yields a relatively low score of 125.9 as shown in Table 10. Had the field not been so heavily influenced by the rise of LLMs, such limitations in small model architectures might have been identified and addressed earlier.
@@
We adopt the LLaVA architecture and inject visual features into the self-attention inputs, as shown in Figure 11 (a). We show that this simple setup surprisingly delivers stronger performance (i.e., Note 1 in Section 3) and contributes to demonstrating that small models can be viable for practical applications. We hope this encourages a shift toward reducing dependence on LLMs in specific tasks like captioning, and sparks further interest in the development of lightweight yet effective models.
E.2 Limitations of Existing Multimodal Large Language Models